# IXPE Observation of the Low-Synchrotron Peaked Blazar S4 0954+65 During An Optical-X-ray Flare

### v1

@@
The X-ray polarization observations made possible with the Imaging X-ray Polarimetry Explorer (IXPE) offer new ways of probing high-energy emission processes in astrophysical jets from blazars. Here we report on the first X-ray polarization observation of the blazar S4 0954+65 in a high optical and X-ray state. During our multi-wavelength campaign on the source, we detected an optical flare whose peak coincided with the peak of an X-ray flare. This optical-X-ray flare most likely took place in a feature moving along the parsec-scale jet, imaged at 43 GHz by the Very Long Baseline Array. The 43 GHz polarization angle of the moving component underwent a rotation near the time of the flare. In the optical band, prior to the IXPE observation, we measured the polarization angle to be aligned with the jet axis. In contrast, during the optical flare the optical polarization angle was perpendicular to the jet axis; after the flare, it reverted to being parallel to the jet axis. Due to the smooth behavior of the optical polarization angle during the flare, we favor shocks as the main acceleration mechanism. We also infer that the ambient magnetic field lines in the jet were parallel to the jet position angle. The average degree of optical polarization during the IXPE observation was (14.34.1)%. Despite the flare, we only detected an upper limit of 14% (at 3 level) on the X-ray polarization degree; although a reasonable assumption on the X-ray polarization angle results in an upper limit of 8.8% (). We model the spectral energy distribution (SED) and spectral polarization distribution (SPD) of S4 0954+65 with leptonic (synchrotron self-Compton) and hadronic (proton and pair synchrotron) models. The constraints we obtain with our combined multi-wavelength polarization observations and SED modeling tentatively disfavor hadronic models for the X-ray emission in S4 0954+65.
@@ 1.1 Blazars
Blazars are active galactic nuclei (AGN) with relativistic jets pointed within a small viewing angle toward an observer. These powerful objects are observed through the entire electromagnetic spectrum from radio to very-high-energy (VHE; 100 GeV to 100 TeV) -rays. Their energy output at almost all wavelengths exhibits rapid, seemingly stochastic variations (e.g., Ulrich et al. 1997). Conventionally, blazars are classified into two subgroups: BL Lac objects (BLLs), with largely featureless optical spectra, and flat spectrum radio quasars (FSRQs), which exhibit strong emission lines. Spectral energy distributions (SEDs) of blazars typically exhibit a double-humped structure. It is generally accepted that the low-energy component of a blazar SED arises due to the synchrotron emission of electrons and positrons radiating in the magnetized jet. Based on the peak frequency of this synchrotron component (), blazars are further subdivided into low-, intermediate-, and high-synchrotron peaked sources (e.g., Abdo et al. 2010): LSP (¡ Hz), ISP (¡¡ Hz), and HSP (¿ Hz). For recent reviews on blazars, we refer the reader to, for example, Böttcher (2019); Blandford et al. (2019); Hovatta & Lindfors (2019).
@@
The origin of the high-energy component of blazar SEDs is still not fully understood. Possibilities include the commonly adopted leptonic models, which explain the high-energy emission as Compton scattering. In these models, the same electrons whose synchrotron emission produces the low-energy SED component upscatter their own synchrotron photons or photons from external sources to higher energies. The former models are referred to as synchrotron self-Compton (SSC; e.g., Jones et al. 1974; Maraschi et al. 1992), while the latter are referred to as external Compton (EC; e.g., Dermer et al. 1992; Sikora et al. 1994). On the other hand, hadronic emission models invoke synchrotron emission by protons, or secondary particles produced in photo-pion cascades, to explain the origin of the high-energy component of blazar SEDs (e.g., Mannheim & Biermann 1989). Hadronic emission models are typically less favored than leptonic ones (e.g., Böttcher et al. 2013a) since, in order for them to be efficient, the total energy requirement in protons is often unsustainably large. It can sometimes be even higher than the Eddington luminosity, since a large fraction of protons must reach energies to exceed the photo-pion energy threshold (e.g., Boettcher et al. 2022). Additionally, hadronic models require rather high magnetic field strengths (e.g., Liodakis & Petropoulou 2020). Nevertheless, the recent 3 spatio-temporal association of the blazar TXS 0506+056 with high-energy neutrinos detected by the IceCube Neutrino Observatory (IceCube Collaboration et al. 2018), and the possibility that there may exist a population-based correlation between high-energy neutrinos and blazars (or at least a subpopulation of blazars; e.g., Padovani et al. 2016; Kouch et al. 2024a), suggest that hadronic processes may be important in blazar jets (e.g., Petropoulou et al. 2020).
1.2 Polarization properties of blazar emission
@@
Given the inherent differences in the polarization properties of electron synchrotron, proton synchrotron, and Compton radiation, multi-wavelength (MWL) polarization measurements are crucial in distinguishing between different particle acceleration and emission models present in blazar jets (e.g., Marscher et al. 2008). Linear polarization is typically represented via two quantities derived from the Stokes parameters (Q, U, and I)111Another Stokes parameter V measures circular polarization, which is intrinsically small in blazars (; Liodakis et al. 2022a, 2023). It is usually ignored and not measured by IXPE at all. When including V, the polarization degree is ).: the polarization degree () and the polarization angle (). The former measures the net fraction of the incoming radiation that is linearly polarized, while the latter is the direction of the linear polarization electric vector on the sky. In blazars, the polarization degree of synchrotron radiation () ranges from zero to as high as 50% (e.g., Mead et al. 1990) and generally indicates how ordered the magnetic field lines are in the emission region. Both electron and proton synchrotron radiation are expected to yield a high degree of polarization if the magnetic field is not very disordered. One way to distinguish them is by measuring the variability of , since that of proton synchrotron is expected to be more stable due to the longer proton cooling times compared to electrons and positrons (Zhang et al. 2016).
@@
Such contrasts between the polarization properties of the various emission models make contemporaneous MWL polarization measurements powerful diagnostic tools for constraining emission and acceleration models, as well as magnetic field strength and morphology in blazar jets, especially when probing higher-energy bands such as X-rays and -rays (e.g., Zhang & Böttcher 2013; Zhang et al. 2019). Notably, in HSP blazars the peak of the low-energy (synchrotron) component falls close to the X-ray regime, hence HSP sources are expected to show a high degree of polarization in the X-ray band. This expectation was confirmed with the advent of the Imaging X-ray Polarimetry Explorer (IXPE, Weisskopf et al. 2022), which has opened a new window of polarimetric measurements to the X-ray universe (2–8 keV; e.g., Ehlert et al. 2022); the highest detected thus far is 31%, in the HSP blazar PKS 2155304 (Kouch et al. 2024b). Altogether, MWL campaigns of several HSP sources conducted contemporaneously by the IXPE collaboration have revealed a general pattern: while there exists a lack of correlation between the MWL polarization properties, the polarization degree in the X-ray band is on average greater than that in the optical band, which is itself on average greater than that in the radio band (i.e., ). This favors acceleration at a shock front as the most likely scenario of particle energization in the jets of HSP sources, with energy stratification in the region beyond the front where the particles lose energy to radiation and adiabatic expansion (e.g., Marscher & Gear 1985; Liodakis et al. 2022b; Di Gesu et al. 2022a; Middei et al. 2023b; Di Gesu et al. 2023a; Ehlert et al. 2023; Kim et al. 2024; Kouch et al. 2024b; Chen et al. 2024). On the other hand, in the case of one HSP source, 1ES 1959+650, which exhibited greater variability than could be resolved by IXPE, the aforementioned trend was not detected (Errando et al. 2024).
@@
In LSP and ISP sources, the X-ray regime generally probes the high-energy component or a mix of the low-energy and high-energy components of the blazar emission. Thus, a sizable (i.e., ) could suggest that hadronic processes play a key role in the high-energy emission of blazars, or that the majority of the high-energy emission can be explained by the anisotropic-EC scenario. On the other hand, a low value of (i.e., ) in LSPs and ISPs would suggest that their high-energy component is dominated by SSC. Likewise, would strongly favor an EC or SSC dominated emission process where the relativistically moving plasma does not contain relatively cold electrons. IXPE observations of the blazar BL Lacertae (an LSP-ISP source) in LSP-like states measured (at 99% confidence level) and demonstrated that (Middei et al. 2023a). Meanwhile, during an ISP-like state when the X-ray photons were coming from the low-energy (synchrotron) component, an IXPE observation of BL Lacertae resulted in in the limited energy range of 2–4 keV (Peirson et al. 2023). IXPE observations of other LSP-ISP sources (three LSPs and one ISP) have only resulted in upper limits of to (at 99% confidence levels; Marshall et al. 2024). Therefore, while these results are not conclusive enough to completely exclude the presence of some hadronic processes in blazar jets, they strongly disfavor purely hadronic (e.g., pure proton synchrotron) and anisotropic-EC scenarios.
1.3 S4 0954+65: a flaring LSP blazar
@@
In this paper we present the results of a MWL campaign of the IXPE collaboration on the LSP blazar S4 0954+65 during an optical-X-ray flaring episode. S4 0954+65 has J2000 sky coordinates and and a redshift of (Lawrence et al. 1986; Becerra González et al. 2021). It is generally classified as a BLL due to its rather weak emission lines (Ghisellini et al. 2011). Nevertheless, the features of its SED somewhat resemble those of an FSRQ, which suggests that S4 0954+65 may be a transitional object between the two classes (Becerra González et al. 2021). S4 0954+65 is a known emitter of GeV and VHE -rays (Mukherjee et al. 1995; Krauss 2014; MAGIC Collaboration et al. 2018). Photometric observations of this source have found long-term variability in both the radio and optical bands (with the former lagging behind the latter by 3 weeks), which was interpreted to arise from dynamical rotation of the jet that gradually alters the beaming angle with respect to our line of sight (e.g., Raiteri et al. 1999, 2021, 2023). Timescale analyses by Sharma et al. (2024) of the 15-year-long Fermi-LAT light curves of S4 0954+65 have led to speculation that the data indicate disc-jet coupling. Furthermore, both photometric and polarimetric observations of S4 0954+65 have shown rapid variability in both the radio and optical bands (e.g., Gabuzda et al. 2000; Marchili et al. 2012; Bachev 2015; Raiteri et al. 2021; Pandey et al. 2023; Raiteri et al. 2023).
@@
Morozova et al. (2014) observed S4 0954+65 as part of a MWL campaign in 2011. They performed optical photometry and polarimetry measurements, as well as Very Long Baseline Array (VLBA) imaging of the source, and utilized Fermi -ray flux density measurements in their study. Between January and March, the optical brightness of S4 0954+65 increased by 3 mag. Between 2011 March-April they observed several optical--ray flares, accompanied by smooth rotations of the optical polarization angle (). During the night of March 9, S4 0954+65 exhibited an extremely fast optical--ray flare (its optical brightness increased by 0.7 mag within 7 h); this occurred along with a smooth rotation in . This was likely associated with the ejection of an apparently superluminal (19) knot from the compact “core” of S4 0954+65. Furthermore, they analyzed other optical flares between 2008 and 2012, finding similar connections between optical variability and the ejection of superluminal knots from the core. Notably, they found that S4 0954+65 generally shows pronounced optical variability (both in flux and polarization properties). Interestingly, they also determined that S4 0954+65 tends to eject around three superluminal knots per year, similar to the rate of optical flares.
@@
Raiteri et al. (2023) conducted a high-cadence photometric and polarimetric monitoring campaign on S4 0954+65 in the optical band during another major flaring episode (April-May 2022). They detected an extremely short variability timescale of 17 minutes, and used it to estimate the size of the emission region to be on the order of pc. They linked such fast, bluer-when-brighter variability patterns to SSC processes present in the jet of the source. Additionally, they found that the variations of the polarization properties in the optical band were more extreme than the radio band, but the average optical polarization angle () and radio polarization angle () remained similar. During this flaring period, on 2022 April 13, Pandey et al. (2023) detected a polarization angle rotation within a 3 hour period in the optical band. They also measured a change as large as 10% in the optical polarization degree of the source within one day. Additionally, they observed that, from mid-March to the end of April 2022, S4 0954+65 reached minimum and maximum values of of 3% and 39%, respectively. In this period, the range of was to . Notably, no correlation between the optical flux density and variations was found (Raiteri et al. 2021; Pandey et al. 2023).
Milliarcsecond-scale (mas) very-long-baseline interferometry (VLBI) maps of S4 0954+65 in the 5 and 8 GHz bands in the 1990s revealed a bright, compact “core” plus a jet pointing west-northwest, bending with distance to point almost directly westward (i.e., )222Position angles are measured from north through east. 3-5 mas from the core (Gabuzda et al., 2000). 15 GHz VLBI maps of S4 0954+65, as part of the Monitoring Of Jets in Active galactic nuclei with VLBA Experiments (MOJAVE333https://www.cv.nrao.edu/MOJAVE/sourcepages/0954+658.shtml) program between the years 1995 and 2013, revealed the jet to point in the general northwest direction on mas scales (; Hovatta et al. 2014; Lister et al. 2018). In 43 GHz VLBA images on sub-mas scales between 2010 and 2018, the jet was in the north-northwest direction (; Jorstad et al. 2017; MAGIC Collaboration et al. 2018; Weaver et al. 2022). The most recent 43 GHz VLBA maps of the source444https://www.bu.edu/blazars/VLBA_GLAST/0954.html suggest that the parsec-scale jet of S4 0954+65 still pointed toward the general north-northwest direction through 2023 (for more details see §3.3).
We describe the X-ray and the accompanying contemporaneous MWL observations of S4 0954+65 in §2 and §3, respectively. We explain the process used to theoretically model the emission of the source in §4. Finally, we summarize the results and discuss them in the context of leptonic and hadronic models in §5.
2 X-ray observations & analysis methods

### v2

@@
The X-ray polarization observations, made possible with the Imaging X-ray Polarimetry Explorer (IXPE), offer new ways of probing high-energy emission processes in astrophysical jets from blazars. Here, we report the first X-ray polarization observation of the blazar S4 0954+65 in a high optical and X-ray state. During our multi-wavelength (MWL) campaign of the source, we detected an optical flare whose peak coincided with the peak of an X-ray flare. This optical-X-ray flare most likely took place in a feature moving along the parsec-scale jet, imaged at 43 GHz by the Very Long Baseline Array (VLBA). The 43 GHz polarization angle of the moving component underwent a rotation near the time of the flare. In the optical band, prior to the IXPE observation, we measured the polarization angle to be aligned with the jet axis. In contrast, during the optical flare, the optical polarization angle was perpendicular to the jet axis; after the flare, it reverted to being parallel to the jet axis. Due to the smooth behavior of the optical polarization angle during the flare, we favor shocks as the main acceleration mechanism. We also infer that the ambient magnetic field lines in the jet were parallel to the jet position angle. The average degree of optical polarization during the IXPE observation was (14.34.1)%. Despite the flare, we only detected an upper limit of 14% (at 3 level) on the X-ray polarization degree; however, a reasonable assumption on the X-ray polarization angle results in an upper limit of 8.8% (). We modeled the spectral energy distribution (SED) and spectral polarization distribution (SPD) of S4 0954+65 with leptonic (synchrotron self-Compton) and hadronic (proton and pair synchrotron) models. Our combined MWL polarization observations and SED modeling tentatively disfavor the use of hadronic models for the X-ray emission in S4 0954+65.
@@
Blazars are active galactic nuclei (AGNs) with relativistic jets aligned within a small viewing angle toward an observer. These powerful objects are observed through the entire electromagnetic spectrum from radio to very-high-energy (VHE; 100 GeV to 100 TeV) -rays. Their energy output at almost all wavelengths exhibits rapid, seemingly stochastic variations (e.g., Ulrich et al. 1997). Conventionally, blazars are classified into two subgroups: BL Lac objects (BLLs), with largely featureless optical spectra, and flat spectrum radio quasars (FSRQs), which exhibit strong emission lines. Spectral energy distributions (SEDs) of blazars typically exhibit a double-humped structure. It is generally accepted that the low-energy component of a blazar SED arises due to the synchrotron emission of electrons and positrons radiating in the magnetized jet. Based on the peak frequency of this synchrotron component (), blazars have been further subdivided into low-, intermediate-, and high-synchrotron peaked sources (e.g., Abdo et al. 2010): LSP (¡ Hz), ISP (¡¡ Hz), and HSP (¿ Hz). For recent reviews of blazar studies, we refer to Böttcher (2019), Blandford et al. (2019), and Hovatta & Lindfors (2019).
@@
The origin of the high-energy component of blazar SEDs is still not fully understood. Possibilities include the commonly adopted leptonic models, which explain the high-energy emission as Compton scattering. In these models, the same electrons whose synchrotron emission produces the low-energy SED component upscatter their own synchrotron photons or photons from external sources to higher energies. The former models are referred to as synchrotron self-Compton (SSC; e.g., Jones et al. 1974; Maraschi et al. 1992), while the latter are referred to as external Compton (EC; e.g., Dermer et al. 1992; Sikora et al. 1994). On the other hand, hadronic emission models invoke synchrotron emission by protons or secondary particles produced in photo-pion cascades, to explain the origin of the high-energy component of blazar SEDs (e.g., Mannheim & Biermann 1989). Hadronic emission models are typically less favored than leptonic ones (e.g., Böttcher et al. 2013), since, for them to be efficient, the total energy requirement in protons is often unsustainably large. It can sometimes be even higher than the Eddington luminosity, since a large fraction of protons must reach energies to exceed the photo-pion energy threshold (e.g., Boettcher et al. 2022). Additionally, hadronic models require rather high magnetic field strengths (10–300 G, e.g., Liodakis & Petropoulou 2020). Nevertheless, the recent 3 spatio-temporal association of the blazar TXS 0506+056 with high-energy neutrinos detected by the IceCube Neutrino Observatory (IceCube Collaboration et al. 2018), coupled with the possibility that there may be a population-based correlation between high-energy neutrinos and blazars (or at least a subpopulation of blazars; e.g., Padovani et al. 2016; Kouch et al. 2024a), suggests that hadronic processes may be important in blazar jets (e.g., Petropoulou et al. 2020).
@@
Given the inherent differences in the polarization properties of electron synchrotron, proton synchrotron, and Compton radiation, multi-wavelength (MWL) polarization measurements are crucial for distinguishing between different particle acceleration and emission models for blazar jets (e.g., Marscher et al. 2008). Linear polarization is typically represented via two quantities derived from the Stokes parameters (Q, U, and I)111Another Stokes parameter V measures circular polarization, which is intrinsically small in blazars (; Liodakis et al. 2022a, 2023). It is usually ignored and not measured by IXPE at all. When including V, the polarization degree is ).: the polarization degree () and the polarization angle (). The former measures the net fraction of the incoming radiation that is linearly polarized, while the latter is the direction of the linear polarization electric vector on the sky. In blazars, the polarization degree of synchrotron radiation () ranges from zero to as high as 50% (e.g., Mead et al. 1990) and generally indicates how ordered the magnetic field lines are in the emission region. Both electron and proton synchrotron radiation are expected to yield a high degree of polarization if the magnetic field is not very disordered. One way to distinguish them is by measuring the variability of , since that of the proton synchrotron is expected to be more stable due to the longer proton cooling times compared to electrons and positrons (Zhang et al. 2016).
@@
Such contrasts between the polarization properties of the various emission models mean that contemporaneous MWL polarization measurements can serve as powerful diagnostic tools for constraining emission and acceleration models, as well as the magnetic field strength and morphology in blazar jets, especially when probing higher-energy bands, such as X-rays and -rays (e.g., Zhang & Böttcher 2013; Zhang et al. 2019). Notably, in HSP blazars the peak of the low-energy (synchrotron) component falls close to the X-ray regime; hence, HSP sources are expected to show a high degree of polarization in the X-ray band. This expectation was confirmed with the advent of the Imaging X-ray Polarimetry Explorer (IXPE, Weisskopf et al. 2022), which has opened a new window of polarimetric measurements to the X-ray universe (2–8 keV; e.g., Ehlert et al. 2022); the highest detected thus far is 31%, in the HSP blazar PKS 2155304 (Kouch et al. 2024b). Altogether, MWL campaigns of several HSP sources conducted contemporaneously by the IXPE collaboration have revealed a general pattern: while there is a lack of correlation between the MWL polarization properties, the polarization degree in the X-ray band is on average greater than that in the optical band, which is itself (on average) greater than that in the radio band (i.e., ). This favors acceleration at a shock front as the most likely scenario of particle energization in the jets of HSP sources, with energy stratification in the region beyond the front where the particles lose energy to radiation and adiabatic expansion (e.g., Marscher & Gear 1985; Liodakis et al. 2022b; Di Gesu et al. 2022a; Middei et al. 2023b; Di Gesu et al. 2023; Ehlert et al. 2023; Kim et al. 2024; Kouch et al. 2024b; Chen et al. 2024). On the other hand, in the case of one HSP source, 1ES 1959+650, it exhibited greater variability than could be resolved by IXPE and the aforementioned trend was not detected (Errando et al. 2024).
@@
In LSP and ISP sources, the X-ray regime generally probes the high-energy component or a mix of the low-energy and high-energy components of the blazar emission. Thus, a sizable (i.e., ) could suggest that hadronic processes play a key role in the high-energy emission of blazars or that the majority of the high-energy emission can be explained by the anisotropic-EC scenario. On the other hand, a low value of (i.e., ) in LSPs and ISPs would suggest that their high-energy component is dominated by SSC. Likewise, would strongly favor an EC or SSC dominated emission process where the relativistically moving plasma does not contain relatively cold electrons. The IXPE observations of the blazar BL Lacertae (an LSP-ISP source) in LSP-like states measured (at 99% confidence level) and demonstrated that (Middei et al. 2023a). Meanwhile, during an ISP-like state when the X-ray photons were coming from the low-energy (synchrotron) component, an IXPE observation of BL Lacertae resulted in in the limited energy range of 2–4 keV (Peirson et al. 2023). The IXPE observations of other LSP-ISP sources (three LSPs and one ISP) have only resulted in upper limits of to (at 99% confidence levels; Marshall et al. 2024). Therefore, while these results are not conclusive enough to completely exclude the presence of some hadronic processes in blazar jets, they strongly disfavor purely hadronic (e.g., pure proton synchrotron) and anisotropic-EC scenarios.
@@
In this paper, we present the results of a MWL campaign of the IXPE collaboration on the LSP blazar S4 0954+65 during an optical-X-ray flaring episode. S4 0954+65 has J2000 sky coordinates of and at a redshift of (Lawrence et al. 1986; Becerra González et al. 2021). It is known to frequently exhibit high polarization and flaring behavior in the optical band. It was identified among the prime targets for IXPE to distinguish between leptonic and hadronic scenarios (Liodakis et al. 2019). S4 0954+65 is generally classified as a BLL due to its rather weak emission lines (Ghisellini et al. 2011). Nevertheless, the features of its SED somewhat resemble those of an FSRQ, which suggests that S4 0954+65 may be a transitional object between the two classes (Becerra González et al. 2021). S4 0954+65 is a known emitter of GeV and VHE -rays (Mukherjee et al. 1995; Krauss 2014; MAGIC Collaboration et al. 2018). Photometric observations of this source have found long-term variability in both the radio and optical bands (with the former lagging behind the latter by 3 weeks), which was interpreted to arise from dynamical rotation of the jet that gradually alters the beaming angle with respect to our line of sight (e.g., Raiteri et al. 1999, 2021, 2023). Timescale analyses by Sharma et al. (2024) of the 15-year-long Fermi-LAT light curves of S4 0954+65 have led to the speculation that the data indicate disc-jet coupling. Furthermore, both photometric and polarimetric observations of S4 0954+65 have shown rapid variability in both the radio and optical bands (e.g., Gabuzda et al. 2000; Marchili et al. 2012; Bachev 2015; Raiteri et al. 2021; Pandey et al. 2023; Raiteri et al. 2023).
@@
Morozova et al. (2014) observed S4 0954+65 as part of a MWL campaign in 2011. They performed optical photometry and polarimetry measurements, as well as Very Long Baseline Array (VLBA) imaging of the source, and utilized Fermi -ray flux density measurements in their study. Between January and March, the optical brightness of S4 0954+65 increased by 3 mag. Between March and April 2011, they observed several optical--ray flares, accompanied by smooth rotations of the optical polarization angle (). On the night of March 9, S4 0954+65 exhibited an extremely fast optical--ray flare (its optical brightness increased by 0.7 mag within 7 h); this occurred along with a smooth rotation in . This was likely associated with the ejection of an apparently superluminal (19) knot from the compact “core” of S4 0954+65. Furthermore, they analyzed other optical flares between 2008 and 2012, finding similar connections between optical variability and the ejection of superluminal knots from the core. Notably, they found that S4 0954+65 generally shows pronounced optical variability (both in flux and polarization properties). Interestingly, they also determined that S4 0954+65 tends to eject around three superluminal knots per year, similar to the rate of optical flares.
@@
Raiteri et al. (2023) conducted a high-cadence photometric and polarimetric monitoring campaign on S4 0954+65 in the optical band during another major flaring episode (April-May 2022). They detected an extremely short variability timescale of 17 minutes, and used it to estimate the size of the emission region to be on the order of pc. They linked such fast, bluer-when-brighter variability patterns to SSC processes present in the jet of the source. Additionally, they found that the variations of the polarization properties in the optical band were more extreme than the radio band, but the average optical polarization angle () and radio polarization angle () remained similar. During this flaring period, on April 13, 2022 , Pandey et al. (2023) detected a polarization angle rotation within a 3 hour period in the optical band. They also measured a change as large as 10% in the optical polarization degree of the source within one day. Additionally, they observed that, from mid-March to the end of April 2022, S4 0954+65 reached minimum and maximum values of of 3% and 39%, respectively. In this period, the range of was to . Notably, no correlation between the optical flux density and variations was found (Raiteri et al. 2021; Pandey et al. 2023).
Milliarcsecond-scale (mas) very-long-baseline interferometry (VLBI) maps of S4 0954+65 in the 5 and 8 GHz bands in the 1990s revealed a bright, compact “core,” plus a jet pointing west-northwest, bending with distance to point almost directly westward (i.e., )222Position angles are measured from north through east. 3-5 mas from the core (Gabuzda et al., 2000). 15 GHz VLBI maps of S4 0954+65, as part of the Monitoring Of Jets in Active galactic nuclei with VLBA Experiments (MOJAVE333https://www.cv.nrao.edu/MOJAVE/sourcepages/0954+658.shtml) program between the years 1995 and 2013, revealed the jet to point in the general northwest direction on mas scales (; Hovatta et al. 2014; Lister et al. 2018). In 43 GHz VLBA images on sub-mas scales between 2010 and 2018, the jet was in the north-northwest direction (; Jorstad et al. 2017; MAGIC Collaboration et al. 2018; Weaver et al. 2022). The most recent 43 GHz VLBA maps of the source444https://www.bu.edu/blazars/VLBA_GLAST/0954.html suggest that the parsec-scale jet of S4 0954+65 still pointed toward the general north-northwest direction throughout 2023 (for more details, see Sects. 3.3).
We describe the X-ray and the accompanying contemporaneous MWL observations of S4 0954+65 in Sects. 2 and 3, respectively. We explain the process used to theoretically model the emission of the source in Sect. 4. Finally, we summarize the results and discuss them in the context of leptonic and hadronic models in Sect. 5.
2 X-ray observations and analysis methods